Rohan's grandmother is five times his age.After 10 years from now ,his grandmother will be three times his age .How old are they at present?
Answer by algebrahouse.com(1659)

x = Rohan's age now
5x = grandmother's age now

x + 10 = Rohan's age after 10 years {added 10 onto current age}
5x + 10 = grandmother's age after 10 years {added 10 onto current age}

5x + 10 = 3(x + 10) {after 10 years, grandmother will be 3 times Rohan's age}
5x + 10 = 3x + 30 {used distributive property}
2x = 20 {subtracted 3x and 10 from each side}
x = 10 {divided each side by 2}
5x = 50 {substituted 10, in for x, into 5x}

Rohan is 10
Grandmother is 50

To ask a question, visit: www.algebrahouse.com